ACCURATE REPS FOR THE SIXTH-ROW MAIN GROUP ELEMENTS

S. A. WILDMAN, G. A. DILABIO, and \underline{P. A. CHRISTIANSEN}, Department of Chemistry, Clarkson University, Potsdam, NY 13699-5810.

Errors in predicted bond lengths in effective potential calculations involving heavy main-group elements are shown to be the result of the poor representation of the f-shell space. Test calculations for PbO and the hydrides of thallium, lead and bismuth demonstrate that with the inclusion of the 5d, 6s and 6p subshells in the valence space and the proper partitioning of the f-shell valence spinors to form pseudospinors, accurate bond lengths are attainable. The previous reasonable bond lengths from 6s6p potentials appear to be the result of fortuitous error cancellations. New REPs in standard form are provided for Tl, Pb, Bi, At and Rn in electronic format at www.clarkson.edu/∼pac/reps.html.